A UFO going 84 m/s south suddenly experienced a constant acceleration for 97 s during which its displacement was 24,783.5 m north. Determine the final speed of the UFO & the magnitude of the UFO's acceleration.
Answer #1

useng sy = uyt + ayt? → 24783.5 = -84x97 + ax 97² > -ay=[-(84*97) – (24783-5)] ay = + 7 m/s² Acceleration: 7 m/s21 (North in